OpenStudy (anonymous):

A circle is described by the equation (x−12)2+(y−32)2=49. What are the cordinates for the center of the circle and the length of the radius? (-12,-32), 23units (-12,-32), 49units (12,32), 49units (12,32), 23units

OpenStudy (gorv):

equation of circle is (x-x1)^2+(y-y1)^2=r^2 where r = radius and centre is (x1,y1)

OpenStudy (gorv):

try it now and tell what you got ??

OpenStudy (anonymous):

dud i cant lie i never did this before can you show me how to do this

OpenStudy (gorv):

you understood what i wrote above in the equation???

OpenStudy (anonymous):

i think so

OpenStudy (gorv):

if you compare my equation with your equation what is x1 = ???

OpenStudy (anonymous):

2x

OpenStudy (gorv):

x1 is what we subtracted from x

OpenStudy (gorv):

what you subtracted from x in your equation??

OpenStudy (anonymous):

x

OpenStudy (gorv):

x-x = 0

OpenStudy (gorv):

(x-12)2

OpenStudy (gorv):

what is subtracted from x here ??

OpenStudy (anonymous):

12

OpenStudy (gorv):

that is our x1 x1=12

OpenStudy (gorv):

if we compare we can see 12 is subtracted from x soit will be x1

OpenStudy (gorv):

similarly can you tell me y1=??

OpenStudy (anonymous):

y

OpenStudy (gorv):

what is subtracted from y = y1

OpenStudy (gorv):

so y1=??

OpenStudy (anonymous):

1

OpenStudy (gorv):

how ???

OpenStudy (anonymous):

hay im runing out of time this class almost over

OpenStudy (gorv):

in your equation what is subtracted from the y ??

OpenStudy (anonymous):

y1

OpenStudy (gorv):

(x−12)2+(y−32)2=49 i dont see y1 anywhere in this equation ..where do u find y1 in this ??

OpenStudy (anonymous):

32

OpenStudy (gorv):

wow .....so y1=32

OpenStudy (gorv):

x1=12 and y1=32 and our centre is (x1,y1)

OpenStudy (gorv):

plug in the value of x1 and y1

OpenStudy (anonymous):

so it 12,32 49 units

OpenStudy (gorv):

r^2=49 r=7

OpenStudy (anonymous):

so it 23?